OpenStudy (anonymous):

A boat's crew rowed 16 miles downstream with current in 2 hours. The return trip upstream, against current covered the same distance, but took 4 hours. Find the crew's rowing rate in in still water & rate of the current. Crews rowing in still water? The rate of the current?

OpenStudy (anonymous):

LOL, no. college algebra

OpenStudy (anonymous):

Let speed in still water be x. Let current speed be y. Speed upstream is x-y, speed downstream is x+y. Now upstream , x-y=16/4=4. Now downstream x+y=16/2=8. Solve the equations for x and y.

OpenStudy (anonymous):

The crew would be 6?

OpenStudy (anonymous):

..